# Fully homomorphic encryption, outsourced databases, high-performance computing

### v1

@@ Abstract
We propose a new framework for compile-time ciphertext synthesis in fully homomorphic encryption (FHE) systems. Instead of invoking encryption algorithms at runtime, our method synthesizes ciphertexts from precomputed encrypted basis vectors using only homomorphic additions, scalar multiplications, and randomized encryptions of zero. This decouples ciphertext generation from encryption, and enables efficient batch encoding through algebraic reuse. We formalize this technique as a randomized module morphism and prove that it satisfies IND-CPA security. Our proof uses a hybrid game framework that interpolates between encrypted vector instances and reduces adversarial advantage to the indistinguishability of the underlying FHE scheme. This reduction structure captures the security implications of ciphertext basis reuse and structured noise injection. The proposed synthesis primitive supports fast, encryption-free ingestion in outsourced database systems and other high-throughput FHE pipelines. It is compatible with standard FHE APIs and preserves layout semantics for downstream homomorphic operations.
@@ 1.1 Background and Motivation
Fully homomorphic encryption (FHE) Gentry (2009); Fan and Vercauteren (2012); Cheon et al. (2017) has become a foundational tool for secure outsourced computation, especially in the context of untrusted cloud-based data management. One prominent application scenario is outsourced database systems Savvides et al. (2020), where data owners upload encrypted records to untrusted servers while retaining the ability to execute structured queries, such as selections, aggregations, and joins—directly over encrypted data.
While recent advances in FHE have significantly reduced the cost of homomorphic operations Kim et al. (2021), the act of encryption itself remains a major bottleneck. This asymmetry is particularly pronounced in streaming and dynamic database environments, where data ingestion is frequent and continuous, whereas query frequency may be sporadic or lightweight in comparison. In such settings, the performance-critical path is often not homomorphic evaluation, but the initial encryption of data.
To address this, the Rache system T. Tawose et al. (2023) proposed an encryption caching technique in which ciphertexts are precomputed for each individual plaintext value and reused when needed. However, this approach fundamentally assumes that each message is a scalar plaintext. In practice, most FHE applications—including encrypted database systems—rely on batch encoding, where a vector of plaintext values is packed into a single ciphertext using polynomial embeddings. This batching mechanism is essential for scalability and aligns with SIMD-style vectorized processing.
Unfortunately, scalar precomputation schemes such as Rache do not generalize to batched encodings. Attempting to precompute ciphertexts on a per-coordinate basis breaks the structural invariants of the encoded plaintext polynomial, leading to semantically invalid ciphertexts and disrupting the vector semantics needed for homomorphic rotation, slot-wise evaluation, and packed aggregation.
This mismatch raises a fundamental question:
Can we perform ciphertext precomputation at the vector level, while preserving compatibility with FHE’s batched encoding and without performing any encryption at runtime?
@@ 1.2 Proposed Work
In this paper, we propose a new algebraic framework for compile-time ciphertext generation in batched fully homomorphic encryption (FHE) systems. The central idea is to decouple ciphertext generation from encryption by defining a reusable synthesis interface: a randomized module morphism that constructs ciphertexts from precomputed encrypted bases.
We model the plaintext space as a finite $\mathbb{Z}_{t}$-module, where each vector can be expressed as a linear combination of standard basis elements. Prior to deployment, we precompute and cache the encryptions of these basis vectors—each embedded via batching into a polynomial ring such as $R=\mathbb{Z}_{q}[x]/(f(x))$. At runtime, a new plaintext vector $\mathbf{m}=\sum_{i}\alpha_{i}e_{i}$ is encrypted by computing the linear combination $\sum_{i}\alpha_{i}\cdot\mathbf{c}_{i}$, where each $\mathbf{c}_{i}=\mathsf{Enc}(e_{i})$ is a cached ciphertext. This process eliminates all online encryption operations, relying solely on ciphertext-level addition and scalar multiplication.
To ensure semantic security, we augment the synthesized ciphertext with a random masking term drawn from a pool of independently generated encryptions of the zero vector. Each noise ciphertext is sampled offline using fresh randomness, forming a high-entropy ensemble over $\mathsf{Enc}(0)$. This masking strategy maintains IND-CPA indistinguishability while preserving layout compatibility for downstream computation.
The resulting framework enables ciphertext synthesis that is both algebraically sound and cryptographically secure. By defining ciphertext generation as a randomized morphism over the module structure, we obtain a compile-time encryption primitive with provable security guarantees under standard assumptions. Our proof strategy uses a hybrid game reduction that bounds adversarial advantage in terms of the IND-CPA security of the underlying FHE scheme.
Finally, this synthesis-based approach provides practical benefits in system-level deployments. In settings such as outsourced encrypted databases or streaming data pipelines, where high-throughput ingestion is required, our method enables an amortized encryption-free runtime path. It integrates seamlessly with existing FHE APIs, preserves slot semantics, and supports symbolic downstream operations such as rotation, packing, and aggregation.
@@ 1.3 Contributions
  • An algebraic interface for compile-time encryption. We introduce a formal abstraction for ciphertext synthesis based on randomized module morphisms, capturing a reusable and backend-agnostic structure for encrypted vector generation.
  • A concrete realization via basis caching. We instantiate this interface using a precomputed ciphertext basis and randomized zero masking, enabling ciphertext synthesis without any encryption at runtime.
  • Formal security under standard assumptions. We prove that the synthesis procedure satisfies IND-CPA security, using a hybrid game argument that reduces adversarial advantage to the underlying FHE encryption scheme.
  • System-level compatibility and ingestion acceleration. Our framework defines a new primitive for encrypted data ingestion, preserving semantic compatibility with FHE APIs while reducing online encryption costs through compile-time vector-level reuse.
@@ 4.3 Randomization
To ensure semantic security and prevent deterministic ciphertext reuse, we inject a random ciphertext of zero into each synthesized output. Let =$\mathcal{R}=\{\mathbf{r}_{1},\dots,\mathbf{r}_{s}\}$ be a pool of independently generated encryptions of the zero vector:
@@
where each $\mathbf{r}_{j}$ has fresh encryption randomness. During synthesis, we select a random index $j\leftarrow[s]$ and output:
@@
This randomization ensures that multiple encryptions of the same plaintext vector $\mathbf{m}$ yield computationally indistinguishable ciphertexts, satisfying IND-CPA under the assumption that $\mathsf{Enc}$ is semantically secure.
@@ 4.4 Overall Algorithm
We now describe the full pipeline for compile-time encryption via vector synthesis. The process is divided into two stages: (1) an offline precomputation stage in which the ciphertext basis and randomized noise pool are constructed, and (2) a runtime synthesis stage that generates ciphertexts for arbitrary plaintext vectors using only cached ciphertexts and ciphertext-level operations.
Algorithm 1 initializes the ciphertext basis $\mathcal{B}$ by encrypting the standard basis vectors $e_{i}\in\mathbb{Z}_{t}^{d}$ using the public key. These basis vectors are embedded into the polynomial ring and encrypted using the FHE scheme’s batch encoding interface. In addition, a pool $\mathcal{R}$ of randomized encryptions of the all-zero vector is generated. Each $\mathbf{r}_{j}\in\mathcal{R}$ is independently sampled using fresh encryption randomness to ensure semantic variability.
Input : Dimension $d$; number of random noise samples $s$; public key $pk$
Output : Cached basis =$\mathcal{B}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{d}\}$; noise pool =$\mathcal{R}=\{\mathbf{r}_{1},\dots,\mathbf{r}_{s}\}$
@@
2Initialize empty list $\mathcal{B}\leftarrow[\,]$, $\mathcal{R}\leftarrow[\,]$;
@@ 8       Append $\mathbf{c}_{i}$ to $\mathcal{B}$
10for $j=1$ to $s$ do
11       $\mathbf{r}_{j}\leftarrow\mathsf{Enc}_{pk}(\mathbf{0})$ with fresh randomness;
12       Append $\mathbf{r}_{j}$ to $\mathcal{R}$
13
return $(\mathcal{B},\mathcal{R})$
Algorithm 1 PrecomputeBasisAndNoise$(d,s,pk)$
At runtime, Algorithm 2 takes as input a plaintext vector $\mathbf{m}=(m_{1},\dots,m_{d})$ and synthesizes the corresponding ciphertext by computing a linear combination of the precomputed basis ciphertexts:
$\mathbf{c}=\sum_{i=1}^{d}m_{i}\cdot\mathbf{c}_{i}+\mathbf{r},$
where $\mathbf{r}\in\mathcal{R}$ is a randomly chosen noise ciphertext. The final ciphertext is computationally indistinguishable from a freshly encrypted version of $\mathbf{m}$ under IND-CPA security, assuming the underlying encryption scheme satisfies semantic security.
@@ 12       Append $\mathbf{r}_{j}$ to $\mathcal{R}$
     Precomputed basis =$\mathcal{B}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{d}\}$;
     Noise pool =$\mathcal{R}=\{\mathbf{r}_{1},\dots,\mathbf{r}_{s}\}$.
@@ 5       $\mathbf{c}\leftarrow\mathbf{c}+m_{i}\cdot\mathbf{c}_{i}$
7Sample $j\leftarrow\{1,\dots,s\}$ uniformly at random;
@@
Algorithm 2 SynthEnc$(\mathbf{m};\mathcal{B},\mathcal{R})$
@@
The runtime cost of $\mathsf{SynthEnc}$ is linear in the vector dimension $d$, requiring $d$ ciphertext-level scalar multiplications and $d+1$ additions over the ciphertext module $R_{q}^{k}$. Assuming constant-time modular operations—as achieved in NTT-optimized FHE backends—the total runtime per synthesized ciphertext is $\mathcal{O}(d)$. The offline precomputation phase involves $\mathcal{O}(d+s)$ encryption calls, where $d$ is the number of basis vectors and $s$ is the size of the randomized zero pool. Once initialized, the online phase is encryption-free and parallelizable across slots, enabling high-throughput ingestion without repeated use of cryptographic primitives.
@@ 5 Correctness and Noise Analysis
In this section, we analyze the correctness and noise behavior of the proposed ciphertext synthesis procedure under standard fully homomorphic encryption (FHE) semantics. Our focus is to verify that synthesized ciphertexts decrypt correctly to their intended plaintext vectors and remain compatible with subsequent homomorphic operations.
5.1 Decryption
Let $\mathbf{c}\in R_{q}^{k}$ be a ciphertext synthesized via Algorithm 2, i.e.,
$\mathbf{c}=\sum_{i=1}^{d}m_{i}\cdot\mathsf{Enc}(e_{i})+\mathsf{Enc}(0),$
where each $\mathsf{Enc}(e_{i})=(c_{i,0},c_{i,1})$ is a ciphertext encrypting a basis vector $e_{i}\in\mathbb{Z}_{t}^{d}$ under standard polynomial encoding.
Decryption proceeds using the secret key $sk=s(x)$:
$\mathsf{Dec}_{sk}(\mathbf{c})=\left(\sum_{i=1}^{d}m_{i}\cdot e_{i}\right)+e(x)=\mathbf{m}+e(x)\in R_{q},$
where $e(x)$ is the aggregated noise term arising from the homomorphic additions and scalar multiplications.
Correct decryption holds as long as:
$\|e(x)\|_{\infty}<\frac{q}{2t},$
ensuring that reduction modulo $t$ eliminates noise and yields the correct plaintext vector.

### v2

@@ Abstract
We propose a framework for compile-time ciphertext synthesis in fully homomorphic encryption (FHE) systems, where ciphertexts are constructed from precomputed encrypted basis vectors combined with a runtime-scaled encryption of zero. This design eliminates online encryption and instead relies solely on ciphertext-level additions and scalar multiplications, enabling efficient data ingestion and algebraic reuse. We formalize the method as a randomized $\mathbb{Z}_{t}$-module morphism and prove that it satisfies IND-CPA security under standard assumptions. The proof uses a hybrid game reduction, showing that adversarial advantage in distinguishing synthesized ciphertexts is negligible if the underlying FHE scheme is IND-CPA secure. Unlike prior designs that require a pool of random encryptions of zero, our construction achieves equivalent security using a single zero ciphertext multiplied by a fresh scalar at runtime, reducing memory overhead while preserving ciphertext randomness. The resulting primitive supports efficient integration with standard FHE APIs and maintains compatibility with batching, rotation, and aggregation, making it well-suited for encrypted databases, streaming pipelines, and secure compiler backends.
@@ 1.1 Background and Motivation
Fully homomorphic encryption (FHE) Gentry (2009); Fan and Vercauteren (2012); Cheon et al. (2017) enables computation over encrypted data without decryption, offering a mathematically rigorous foundation for secure delegated computation. Among its most impactful application domains is outsourced data processing, where sensitive data is stored and queried in untrusted cloud infrastructures. A canonical example is encrypted database-as-a-service (DBaaS) Savvides et al. (2020), in which the data owner uploads encrypted tables to the cloud and delegates query execution—such as selection, aggregation, joins, or even machine learning inference—while preserving data confidentiality.
While the last decade has witnessed dramatic improvements in the efficiency of homomorphic arithmetic SEAL ; Kim et al. (2021), a critical asymmetry persists in system-level FHE deployments: the overhead of encryption remains disproportionately high compared to that of homomorphic evaluation. This is particularly acute in dynamic and streaming workloads, such as real-time monitoring, log ingestion, and encrypted sensor pipelines, where large volumes of new data must be continuously encrypted and injected into the system. In such contexts, the primary bottleneck shifts from query latency to data ingestion throughput, with encryption emerging as the dominant cost in the trusted domain.
Several systems have attempted to address this bottleneck through precomputation strategies. Notably, the Rache system T. Tawose et al. (2023) introduced a form of encryption caching, wherein the ciphertexts for all possible scalar plaintexts are pre-encrypted and stored in a lookup table. This approach enables constant-time retrieval of ciphertexts without executing encryption algorithms at runtime. However, Rache and similar scalar-based schemes are intrinsically tied to unbatched representations, treating each message as a standalone integer or boolean. In contrast, real-world FHE applications overwhelmingly employ batch encoding, which maps entire vectors of plaintext values into the coefficient structure of a single polynomial in $R=\mathbb{Z}_{q}[x]/(f(x))$. This encoding supports parallel SIMD-style execution across slots, and is central to the performance of encrypted matrix operations, packed aggregation, and neural inference.
Unfortunately, scalar-level caching strategies are incompatible with these batched encodings. Naively precomputing encrypted vectors on a per-slot basis does not preserve the algebraic structure imposed by the batching transformation, and leads to ciphertexts that violate semantic layout invariants. These malformed ciphertexts cannot be consumed by standard FHE operations such as $\mathsf{EvalRotate}$ or $\mathsf{EvalSum}$, and break homomorphic slot alignment required by higher-level algorithms. More fundamentally, precomputing all possible batched vectors is combinatorially infeasible due to the exponential growth in vector space size.
This tension gives rise to a central open question for compile-time FHE system design:
Is it possible to perform ciphertext precomputation at the vector level, while preserving compatibility with FHE’s batching interface and without invoking encryption at runtime?
In this paper, we answer this question in the affirmative. We present a principled framework for ciphertext synthesis over batched encodings, in which encrypted vectors are constructed via compile-time algebraic composition of a precomputed ciphertext basis and a lightweight randomized masking term. This approach bridges the gap between symbolic data generation and cryptographic soundness, and lays the foundation for high-throughput ingestion pipelines in encrypted database systems and beyond.
@@ 1.2 Proposed Work
We propose a new algebraic abstraction for compile-time ciphertext synthesis in batched fully homomorphic encryption (FHE) systems. The core insight is to shift encryption from a runtime operation to a symbolic, algebraic process, where ciphertexts are constructed through linear combinations over a precompiled encrypted basis and masked using a single ciphertext of zero scaled by runtime randomness.
We model the plaintext space as a finite-dimensional $\mathbb{Z}_{t}$-module, and precompute the encryptions of standard basis vectors $\{e_{i}\}_{i=1}^{d}$, each embedded into the polynomial ring $R=\mathbb{Z}_{q}[x]/(f(x))$ via batching. These encrypted basis vectors $\{\mathbf{c}_{i}=\mathsf{Enc}(e_{i})\}$ form a compile-time ciphertext basis $\mathcal{B}$. Given an input vector $\mathbf{m}=\sum_{i}m_{i}e_{i}$, ciphertext generation is performed at runtime by evaluating the linear combination $\sum_{i}m_{i}\cdot\mathbf{c}_{i}$, using ciphertext-level operations only.
To preserve semantic security and avoid determinism, we introduce a randomized masking step via a fixed encryption of the zero vector, denoted $\mathbf{r}_{0}=\mathsf{Enc}(0^{d})$. At runtime, this ciphertext is scaled by a freshly sampled scalar $\rho\leftarrow\mathbb{Z}_{t}$ and added to the synthesized output:
$\textsf{SynthEnc}(\mathbf{m}):=\sum_{i}m_{i}\cdot\mathbf{c}_{i}+\rho\cdot\mathbf{r}_{0}.$
This design maintains ciphertext-level randomness without requiring a pool of zero encryptions, reducing memory footprint and simplifying implementation.
Our construction can be viewed as a randomized module morphism over encrypted $\mathbb{Z}_{t}$-modules. By embedding encryption into compile-time algebra, we obtain a stateless encryption interface that is compatible with downstream homomorphic operations and secure under standard IND-CPA assumptions. The associated security proof is structured as a hybrid game sequence, with reductions to the indistinguishability of the underlying FHE encryption.
Practically, this synthesis interface supports high-throughput ingestion for encrypted data systems, enabling runtime ciphertext generation without invoking cryptographic primitives. The construction is compatible with batching, ciphertext rotations, and slot-wise homomorphic operations, and offers a pathway toward symbolic encrypted compilers and bootstrappable pipelines.
@@ 1.3 Contributions
  • Algebraic abstraction for compile-time encryption. We introduce a novel encryption interface modeled as a randomized module morphism over a precomputed ciphertext basis, enabling deterministic synthesis of ciphertexts with runtime-randomized masking.
  • Minimal-noise masking with provable security. Unlike prior work that requires a pool of random zero encryptions, we show that a single ciphertext of zero suffices when combined with scalar multiplication by a fresh randomizer. This design reduces memory requirements while preserving semantic security.
  • Formal IND-CPA security via hybrid argument. We establish a reduction from the IND-CPA security of synthesized ciphertexts to that of the underlying FHE scheme, using a coordinate-wise hybrid game construction and precise advantage bounding.
  • Practical integration and encryption bypass. Our approach supports encryption-free runtime operation in FHE systems, with immediate application to encrypted databases, ingestion pipelines, and vectorized secure compilers.
@@ 4.3 Randomization
To prevent deterministic reuse and achieve semantic security, we inject a ciphertext encoding of the zero vector into every synthesized encryption. This ensures that identical plaintext vectors yield ciphertexts that are computationally indistinguishable.
In our original formulation, we maintained a pool =$\mathcal{R}=\{\mathbf{r}_{1},\dots,\mathbf{r}_{s}\}$ consisting of independently encrypted zero vectors, each generated as
@@
where each ciphertext $\mathbf{r}_{j}$ contains fresh encryption randomness. During ciphertext synthesis, a random index $j\leftarrow[s]$ is chosen and the encrypted zero vector $\mathbf{r}_{j}$ is added to the linear combination, yielding:
@@
This mechanism ensures IND-CPA security under the assumption that the encryption scheme $\mathsf{Enc}$ is semantically secure.
We now present an optimized design that achieves the same security goal with significantly reduced memory overhead. Instead of maintaining an entire pool $\mathcal{R}$ of encrypted zero vectors, we cache a single ciphertext $\mathbf{r}_{0}\leftarrow\mathsf{Enc}(0^{d})$ and inject fresh randomness at synthesis time via scalar multiplication. Specifically, we sample a fresh scalar $a\leftarrow\mathbb{Z}_{q}$ and compute:
$\textsf{SynthEnc}_{\mathcal{B}}^{\mathbf{r}_{0}}(\mathbf{m}):=\sum_{i=1}^{d}m_{i}\cdot\mathbf{c}_{i}+a\cdot\mathbf{r}_{0}.$
The ciphertext $a\cdot\mathbf{r}_{0}$ remains a valid encryption of the zero vector due to the homomorphic properties of the underlying scheme, and the use of a fresh scalar $a$ guarantees statistical masking across invocations.
This optimization eliminates the need to store $s$ distinct ciphertexts, reducing the memory requirement from $O(s)$ to $O(1)$ while preserving the semantic security of the overall scheme. The randomness space is preserved since the space of scalar multiples of a semantically secure encryption remains indistinguishable under the assumed hardness of the underlying encryption. Consequently, this design achieves IND-CPA security with minimal memory footprint and improved cache reuse.
@@ 4.4 Overall Algorithm
We now describe the full pipeline for compile-time encryption via vector synthesis. The system consists of two phases: an offline phase where the ciphertext basis and reusable zero ciphertext are generated, and an online phase that synthesizes new ciphertexts using only cached values and arithmetic over ciphertexts.
Algorithm 1 initializes the ciphertext basis $\mathcal{B}$ by encrypting the standard basis vectors $e_{i}\in\mathbb{Z}_{t}^{d}$ using the FHE scheme’s batch encoding interface. Each $\mathbf{c}_{i}=\mathsf{Enc}_{pk}(e_{i})$ is a ciphertext encoding the $i$-th unit vector, and the full collection $\mathcal{B}$ enables synthesis of arbitrary plaintexts via linear combination. Additionally, a single ciphertext $\mathbf{r}_{0}=\mathsf{Enc}_{pk}(0^{d})$ is precomputed using fresh encryption randomness. This ciphertext encodes the all-zero vector but will serve as a reusable noise carrier during synthesis.
Input : Dimension $d$; public key $pk$
Output : Cached basis =$\mathcal{B}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{d}\}$; reusable noise ciphertext $\mathbf{r}_{0}$
@@
2Initialize empty list $\mathcal{B}\leftarrow[\,]$;
@@ 8       Append $\mathbf{c}_{i}$ to $\mathcal{B}$
[…]
11
return $(\mathcal{B},\mathbf{r}_{0})$
Algorithm 1 PrecomputeBasisAndNoise$(d,pk)$
Algorithm 2 implements the online phase. Given a plaintext vector $\mathbf{m}=(m_{1},\dots,m_{d})$, the algorithm computes the synthesized ciphertext
$\mathbf{c}=\sum_{i=1}^{d}m_{i}\cdot\mathbf{c}_{i}+a\cdot\mathbf{r}_{0},$
where $a\leftarrow\mathbb{Z}_{q}$ is a runtime-sampled scalar used to mask the zero ciphertext. Because the ciphertext module supports linear operations, the scalar-multiplied zero ciphertext $a\cdot\mathbf{r}_{0}$ acts as a randomized blinding term that guarantees ciphertext uniqueness without requiring fresh encryption calls.
@@
     Cached basis =$\mathcal{B}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{d}\}$;
     Reusable zero ciphertext $\mathbf{r}_{0}=\mathsf{Enc}_{pk}(0^{d})$.
@@ 5       $\mathbf{c}\leftarrow\mathbf{c}+m_{i}\cdot\mathbf{c}_{i}$
7Sample $a\leftarrow\mathbb{Z}_{q}$ uniformly at random;
@@
Algorithm 2 SynthEnc$(\mathbf{m};\mathcal{B},\mathbf{r}_{0})$
@@
The runtime cost of $\mathsf{SynthEnc}$ remains $\mathcal{O}(d)$, requiring $d$ scalar multiplications and $d+1$ additions in the ciphertext domain. Crucially, this procedure performs no encryption during synthesis. The precomputation phase performs $d+1$ encryptions—one for each basis vector and one for the reusable zero vector—but these are amortized across all future invocations. This leads to efficient, parallelizable, and encryption-free synthesis with IND-CPA security guarantees.
@@ 5 Correctness and Noise Analysis
We analyze the correctness and noise behavior of ciphertexts produced via compile-time synthesis. All results are interpreted under standard leveled FHE semantics, where ciphertexts are tuples over a ring $R_{q}$ and correctness is defined by bounded decryption error.
5.1 Decryption Correctness
Let $\mathbf{c}$ be a synthesized ciphertext computed as
$\mathbf{c}=\sum_{i=1}^{d}m_{i}\cdot\mathbf{c}_{i}+a\cdot\mathbf{r}_{0},$
where $\mathbf{c}_{i}=\mathsf{Enc}(e_{i})$ and $\mathbf{r}_{0}=\mathsf{Enc}(0^{d})$ are precomputed ciphertexts, and $a\in\mathbb{Z}_{q}$ is a scalar sampled uniformly at runtime. All ciphertexts are assumed to be in the 2-component format $\mathbf{c}=(c_{0},c_{1})$.
Each term $m_{i}\cdot\mathbf{c}_{i}$ contributes noise linearly scaled by $|m_{i}|$, while the randomization term $a\cdot\mathbf{r}_{0}$ introduces additional noise scaled by $|a|$. Let $e_{i}(x)$ and $e_{r}(x)$ denote the noise polynomials of $\mathbf{c}_{i}$ and $\mathbf{r}_{0}$, respectively. Then, the total decryption error is
$e_{\text{total}}(x)=\sum_{i=1}^{d}m_{i}\cdot e_{i}(x)+a\cdot e_{r}(x).$
Correct decryption requires that the aggregate error remains below the modulus-resolution threshold:
$\|e_{\text{total}}(x)\|_{\infty}<\frac{q}{2t}.$
This condition holds when:
  • Each $m_{i}$ is drawn from a bounded plaintext alphabet (e.g., $|m_{i}|<t/4$),
  • The scalar $a$ is sampled uniformly from a small range (e.g., $\{0,1\}$ or $\mathbb{Z}_{t}$),
  • The noise of each precomputed ciphertext satisfies $\|e_{i}(x)\|_{\infty}\ll q/t$.
These bounds can be enforced offline at synthesis-setup time.